### v1

# Existence and uniqueness of solutions to SDEs with state-dependent variable exponent

Abstract
We study a time-inhomogeneous, nonlinear stochastic differential equation whose drift and diffusion terms involve state-dependent variable exponents. This novel structure introduces flexibility, allowing the system to adapt dynamically to its current state and better capture real-world phenomena compared to classical models such as GBM and CEV. While this flexibility enhances modeling accuracy, it also creates analytical challenges. Using a fixed-point approach, we establish existence and uniqueness under suitable conditions. The model generalizes GBM and CEV while exhibiting distinct behavior for small states.
keywords:
Stochastic process; state-dependent variable exponent; Banach fixed-point theorem.
2020 MSC:
60G07; 60G20; 60G65; 60H15; 60H20; 60H30
journal:
[…]
organization=Faculty of Science and Technology, Applied Mathematics
Athabasca University,city=Athabasca, postcode=T9S 3A3, state=AB, country=Canada
1 Introduction
We study the following time-inhomogeneous nonlinear stochastic differential equation with state-dependent variable exponent functions $p(X(t)),q(X(t))$
$\begin{cases}\begin{array}[]{rll}dX(t)&=\mu(t)X(t)^{p(X(t))}dt+\sigma(t)X(t)^{q(X(t))}dW(t),\,\,t\in[0,T],\\ X_{0}&=x_{0},\quad x_{0}\in(0,\infty),\end{array}\end{cases}$ ($\mathcal{GM}$})
where $W(t)$ is a Brownian motion on the complete probability space $(\Omega,\mathcal{F},\mathbb{P})$; $p(\cdot),q(\cdot)$ are differentiable, variable exponent functions; and $0<\mu,\sigma$ are bounded real-valued deterministic functions.
Stochastic differential equations (SDEs) play a central role in modeling systems influenced by randomness, with applications spanning finance, physics, biology, and engineering [7, 8, 9]. Classical models such as the geometric Brownian motion (GBM) and the constant elasticity of variance (CEV) model have been extensively studied (see e.g., [4, 5]). However, these models often assume fixed exponents in drift and diffusion terms, limiting their ability to capture complex state-dependent dynamics.
The proposed model ($\mathcal{GM}$}), as a novel approach, introduces state-dependent variable exponents $p(\cdot)$ and $q(\cdot)$ in both drift and diffusion terms. Unlike classical models with constant elasticity, this formulation allows the coefficients to respond nonlinearly to the current state. Such flexibility significantly enhances the expressive power of SDEs, enabling better modeling of phenomena where volatility and growth rates vary with the state. However, this nonlinearity creates analytical challenges: standard techniques for proving existence and uniqueness often rely on Lipschitz conditions, which are harder to verify under variable exponent structures. Our work addresses these challenges by introducing a new class of functions, $\mathcal{S}$, for variable exponents to ensure the coefficients satisfy the necessary regularity properties.
The proposed model ($\mathcal{GM}$}) generalizes several classical models. For example:
  • 1.
    If $p(x)=q(x)=1$ and $\mu(t)=\mu,\,\sigma(t)=\sigma$, ($\mathcal{GM}$}) reduces to the GBM. A key feature of hypothesis $(\mathbf{h_{2}})$ is that $p(x)\to 1$ and $q(x)\to 1$ as $x\to\infty$. This implies that our generalized model behaves like GBM for large $x$, but exhibits different dynamics for small $x$.
  • 2.
    If $p(x)=1$ and $q(x)\geq 1$ (excluding $0\leq q(x)<1$) and $\mu(t)=\mu,\,\sigma(t)=\sigma$, the model reduces to the CEV model.
Thus, ($\mathcal{GM}$}) provides a unified framework that encompasses GBM and CEV as special cases while introducing richer dynamics through state-dependent exponents.
In this paper, building on the research findings of [1, 2, 3], we differently study a time-inhomogeneous model, and obtain the existence and uniqueness of solution by exploiting the idea of the contraction mapping principle, also known as the Banach fixed-point theorem.
2 Preliminaries and auxiliary results
We introduce a new class of functions, $\mathcal{S}$, for variable exponents to ensure that the model’s coefficients satisfy Lipschitz and linear growth conditions.
  • ($\mathbf{h_{1}}$)
    $h(\cdot):(0,\infty)\to[1,\infty)$ is a differentiable function satisfying
    $1\leq h^{-}:=\inf_{x>0}h(x)\quad\text{and}\quad\sup_{x>0}h(x):=h^{+}<\infty.$
  • ($\mathbf{h_{2}}$)
    $\lim_{x\to\infty}h(x)=1\quad\text{and}\quad\limsup_{x\to\infty}(h(x)-1)\log(x)<\infty.$
  • ($\mathbf{h_{3}}$)
    There exist real numbers $\delta,M_{0},C_{0}>0$ and $\alpha>0$ with $h^{+}<1+\alpha$ such that
    $|h^{\prime}(x)|\leq M_{0}\mathbf{1}_{(0,\delta]}(x)+C_{0}x^{-(1+\alpha)}\mathbf{1}_{(\delta,\infty)}(x).$
  • $(\mathbf{D})$
    $\mu,\sigma:[0,T]\to(0,\infty)$ are deterministic functions such that
    $0<\mu^{-}:=\inf_{t\in[0,T]}\mu(t),\quad\sup_{t\in[0,T]}\mu(t):=\mu^{+}<\infty.$
    $0<\sigma^{-}:=\inf_{t\in[0,T]}\sigma(t),\quad\sup_{t\in[0,T]}\sigma(t):=\sigma^{+}<\infty.$
We say that the function $h(\cdot)$ belongs to the class $\mathcal{S}$ if it satisfies the hypotheses $(\mathbf{h_{1}})-(\mathbf{h_{3}})$.
Remark 1.
The function $h(x)=1+\frac{0.5}{1+x}$ satisfies the hypotheses $(\mathbf{h_{1}})$-$(\mathbf{h_{3}})$. $(\mathbf{h_{1}})$ and $(\mathbf{h_{2}})$ easily follows. If one lets $\delta=1$, $M_{0}\geq 0.5$, $C_{0}\geq 1$ and $\alpha\geq 1$ then $(\mathbf{h_{3}})$ is also satisfied.
Let $\{\mathcal{F}_{t}\}_{t\geq 0}$ be a filtration denoted by $\mathcal{F}_{t}$. Then $(\Omega,\mathcal{F},\mathcal{F}_{t},\mathbb{P})$ is a filtered probability space satisfying the usual conditions (see, e.g., [8, 9]). A stochastic process $X(t)$, defined for $t\geq 0$, is said to be progressively measurable with respect to a filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$ if, for every $t\in[0,T]$, the mapping
$X:[0,t]\times\Omega\to\mathbb{R},\quad(s,\omega)\mapsto X(s,\omega)$ (1)
is measurable with respect to the product $\sigma$-field $\mathcal{B}([0,t])\otimes\mathcal{F}_{t}$, where $\mathcal{B}([0,t])$ denotes the Borel $\sigma$-field on $[0,t]$. Note that since the stochastic process $X(t)$ given by ($\mathcal{GM}$}) is adapted to $\mathcal{F}_{t}$ and has continuous paths $\mathbb{P}$-almost surely ($\mathbb{P}$-a.s.), it is progressively measurable. In order to show the existence of a solution, we are going to employ the Banach fixed point theorem in the space
$\mathcal{M_{T}}=\left\{X(t):X\text{ is progressively measurable, }\int_{\Omega}\int_{0}^{T}|X(t,\omega)|^{2}dtd\mathbb{P}(\omega)<\infty\right\}$ (2)
with the norm
$\|X\|^{2}_{\mathcal{T}}=\mathbb{E}\left[\int_{0}^{T}|X(t)|^{2}dt\right].$ (3)
The normed space $(\mathcal{M_{T}},\|\cdot\|_{\mathcal{T}})$ is complete since it is a closed subspace of the space $L^{2}([0,T]\times\Omega)$ of functions of two variables $(t,\omega)\in[0,T]\times\Omega$ that are square-integrable with respect to the product measure $d\left(\nu\times\mathbb{P}\right)(t,\omega)=dtd\mathbb{P}(\omega)$, where $\nu$ is Lebesgue measure on $[0,T]$.
Lemma 2.1.
Assume $X(t)\in\mathcal{M_{T}}$ and $h(\cdot)$ satisfies $(\mathbf{h_{1}})$ and $(\mathbf{h_{2}})$. Then $X^{h}(t):=X(t)^{h(X(t))}\in\mathcal{M_{T}}$.
Proof.
It is enough to apply ($\mathbf{h_{2}}$). If $0<x\leq 1$, there exists $0<\varepsilon\leq 1$ such that
$x^{h(x)}\leq\varepsilon^{h^{-}}\leq\varepsilon^{h^{-}}(1+x).$ (4)
If $1<x<\infty$, there exist real numbers $M_{\infty},R_{\infty}>0$ such that
$(h(x)-1)\log(x)\leq M_{\infty},\quad\forall x>R_{\infty},$ (5)
which yields
$e^{(h(x)-1)\log(x)}\leq e^{M_{\infty}}\Rightarrow x^{h(x)}\leq e^{M_{\infty}}(1+x).$ (6)
Therefore, for any $x\in(0,\infty)$
$x^{h(x)}\leq K(1+x),$ (7)
where $K:=\max\{\varepsilon^{h^{-}},e^{M_{\infty}}\}$. Hence, considering all these along with the fact that $X(t)\in\mathcal{M_{T}}$, it follows that
$\mathbb{E}\left[\int_{0}^{T}|X(t)|^{2h(X(t))}dt\right]\leq 2K^{2}(T+1)\mathbb{E}\left[\int_{0}^{T}X^{2}(t)ds\right]=2K^{2}(T+1)\|X\|^{2}_{\mathcal{T}}<\infty.$ (8)
Lemma 2.2.
Assume that $X_{0}=x_{0}>0$, and $p(\cdot),q(\cdot)$ satisfy $(\mathbf{h_{1}})$ and $(\mathbf{h_{3}})$. Then $X(t)$ remains strictly positive in the state space $(0,\infty)$ $\mathbb{P}$-a.s. for all $t\in[0,T]$. In particular, the boundary $x=0$ is natural; that is, starting from $x_{0}$, it is never hit in finite time.
Proof.
We use the local (differential) form of Feller’s non-attainability test [6] to analyze the behavior of the processes $X(t)$ at the boundary of its state space $(0,\infty)$. For a general SDE
$dX(t)=\mu\left(t,X(t)\right)dt+\sigma\left(t,X(t)\right)dW(t)$
if
$\lim_{x\to 0^{+}}\left(\mu(t,x)-\frac{1}{2}\frac{\partial\sigma^{2}}{\partial x}(t,x)\right)\geq 0$ (9)
holds, then the boundary $X(t)=0$ is non-attainable for the process $X(t)$ given that $x_{0}>0$. Now, we apply (9) to ($\mathcal{GM}$}). Then, for $0<x\leq\delta<1$,
$\mathcal{T}(t,x)=\mu(t)x^{p(x)}-\frac{\sigma^{2}(t)}{2}\frac{\partial x^{2q(x)}}{\partial x}\geq\mu^{-}x^{p(x)}-(\sigma^{+})^{2}\left(M_{0}x^{2q(x)}|\log(x)|+x^{2q(x)-1}q^{+}\right).$ (10)
Using the elementary calculus along with the assumptions $(\mathbf{h_{1}})$ and $(\mathbf{h_{3}})$, it follows
$\lim_{x\to 0^{+}}\mathcal{T}(t,x)$ $\geq\lim_{x\to 0^{+}}\left(\mu^{-}x^{p(x)}-(\sigma^{+})^{2}\left(M_{0}x^{2q(x)}|\log(x)|+x^{2q(x)-1}q^{+}\right)\right)=0.$ (11)
Hence $X(t)>0$ $\mathbb{P}$-a.s. for all $t\in[0,T]$, which means ($\mathcal{GM}$}) is well-defined. ∎
3 The main result
Theorem 3.1.
Assume that $p(\cdot),q(\cdot)\in\mathcal{S}$ and condition $(\mathbf{D})$ holds. Also, suppose the initial condition $X_{0}=X(0)=x_{0}$ is $\mathcal{F}_{0}$-measurable, independent of $W(t)$, and has a finite second moment. Then there exists a unique solution $X(t)$ of ($\mathcal{GM}$}) with continuous paths satisfying $\mathbb{E}\left[\int_{0}^{T}|X(t)|^{2}dt\right]<\infty$.
Proof.
Let us fix $T>0$. Let $\Phi$ be the function that maps a process $X(t)$ into a process $\Phi X(t)$ defined by
$\Phi X(t)=x_{0}+\int_{0}^{t}\mu(s)X^{p}(s)ds+\int_{0}^{t}\sigma(s)X^{q}(s)dW(s).$ (12)
In this context, the solution of ($\mathcal{GM}$}) is a process $X(t)$ that satisfies the equation
$\Phi X(t)=X(t),$ (13)
that is, it is a fixed point of $\Phi$. We shall present the proof in four steps.
Step 1. Let $X(t)\in\mathcal{M_{T}}$. Then
$(\Phi X(t))^{2}\leq 3x^{2}_{0}+3(\mu^{+})^{2}\left(\int_{0}^{t}X^{p}(s)ds\right)^{2}+3(\sigma^{+})^{2}\left(\int_{0}^{t}X^{q}(s)dW(s)\right)^{2}.$ (14)
Taking the expectation and integrating with respect to $t$ we obtain
$\int_{0}^{T}\mathbb{E}\left[(\Phi X(s))^{2}\right]dt$ $\leq 3T\mathbb{E}[x^{2}_{0}]+3(\mu^{+})^{2}\int_{0}^{T}\mathbb{E}\left[\left(\int_{0}^{t}X^{p}(s)ds\right)^{2}\right]dt$
$+3(\sigma^{+})^{2}\int_{0}^{T}\mathbb{E}\left[\left(\int_{0}^{t}X^{q}(s)dW(s)\right)^{2}\right]dt.$ (15)
Now, applying the Cauchy-Schwarz inequality to the second term, the Itô isometry to the third term (since $X^{q}(t)\in\mathcal{M_{T}}$) in (3), and using Lemma 2.1, we obtain
$\int_{0}^{T}\mathbb{E}\left[(\Phi X(s))^{2}\right]dt$ $\leq 3T\mathbb{E}[x^{2}_{0}]+3[(\mu^{+})^{2}+(\sigma^{+})^{2}]K^{2}\|X\|^{2}_{\mathcal{T}}\int_{0}^{T}2t(t+1)dt<\infty.$ (16)
Since $(\Phi X(s))^{2}$ is measurable on $0,T]\times\Omega$, by the Fubini-Tonelli theorem, one concludes that $\Phi:\mathcal{M_{T}}\to\mathcal{M_{T}}$. Note that if one follows similar arguments as with the proof of Lemma 3.1 $(i)$ of [2] and consider $(\mathbf{D})$, it is straightforward to show that there exists a constant $L=L(\mu_{+},\sigma_{+})>0$ such that $\forall x,y\in(0,\infty),t\in[0,T]$, it holds:
$|\mu(t)(x^{p(x)}-y^{p(y)})|+|\sigma(t)(x^{q(x)}-y^{q(y)})|\leq L|x-y|,$ (17)
where $\mu_{+}=\max\{\mu^{-},\mu^{+}\}$ and $\sigma_{+}=\max\{\sigma^{-},\sigma^{+}\}$. Let $X(t),Y(t)\in\mathcal{M_{T}}$. Then applying the Cauchy-Schwarz inequality and the Itô isometry alon with (17) gives
$\|\Phi X-\Phi Y\|^{2}_{\mathcal{T}}$ $\leq 2\int_{0}^{T}\left(\mu_{+}^{2}\mathbb{E}\left[\left(\int_{0}^{t}\Delta(s)dt\right)^{2}\right]+\sigma_{+}^{2}\mathbb{E}\left[\left(\int_{0}^{t}\Delta(s)dW(s)\right)^{2}\right]\right)dt$
$\leq c^{2}\|X-Y\|^{2}_{\mathcal{T}},$ (18)
where $\Delta(t):=X(t)-Y(t)$ and $c=\sqrt{2L^{2}T(T\mu_{+}^{2}+\sigma_{+}^{2})}$. Finally, if we choose $0<T_{*}\leq T$ so that $c<1$, it turns out that $\Phi$ is a contraction from $\mathcal{M_{T_{*}}}$ into $\mathcal{M_{T_{*}}}$, i.e.
$\|\Phi X-\Phi Y\|_{\mathcal{T_{*}}}$ $\leq c\|X-Y\|_{\mathcal{T_{*}}}.$ (19)
Step 2. Let $X_{n}(t)$ be a sequence in $\mathcal{M_{T_{*}}}$. Define the sequence
$X_{n+1}(t)=\Phi X_{n}(t)\quad\text{with}\quad X_{1}(t)=\Phi X_{0}(t).$
Then consecutive approximations and induction lead to
$\|X_{n+1}-X_{n}\|_{\mathcal{T_{*}}}\leq c^{n}\|X_{1}-X_{0}\|_{\mathcal{T_{*}}}.$ (20)
Thus, for $n<m$
$\|X_{m}-X_{n}\|_{\mathcal{T_{*}}}\leq(c^{m-1}+c^{m-2}+\cdots+c^{n})\|X_{1}-X_{0}\|_{\mathcal{T_{*}}}.$ (21)
Since $c<1$, the sum of geometric series $\sum_{k=n}^{m-1}c^{k}$ tends to 0 as $n,m\to\infty$. Thus, $X_{n}$ is a Cauchy sequence in $\mathcal{M_{T_{*}}}$. Since $\mathcal{M_{T_{*}}}$ is complete and $\Phi$ is continuous, passing to the limit in $X_{n+1}(t)=\Phi X_{n}(t)$ gives $X(t)=\Phi X(t)$. Thus, $X(t)$ is a solution of ($\mathcal{GM}$}) in $\mathcal{M_{T_{*}}}$. This solution is unique. Argue otherwise and assume that $\Phi$ has another fixed point, $Y(t)\in\mathcal{M_{T_{*}}}$ such that $X(t)\neq Y(t)$ $\mathbb{P}$-a.s. Then
$\|\Phi X-\Phi Y\|_{\mathcal{T_{*}}}$ $<\|X-Y\|_{\mathcal{T_{*}}}=\|\Phi X-\Phi Y\|_{\mathcal{T_{*}}},$ (22)
which is a contradiction.
Step 3. We already solved the equation on $0,T_{*}]$ with initial condition $X(0)$. Now we want to ensure that the solution can be continued to $T_{*},2T_{*}]$, then $2T_{*},3T_{*}]$, and so on, until we cover $[0,T]$. In doing so, by applying the same arguments as before, we obtain for $t\leq T_{*}$
$\mathbb{E}\left[X^{2}(t)\right]$ $\leq 3\mathbb{E}[x^{2}_{0}]+3[2(T+1)K^{2}(\mu^{2}_{+}+\sigma^{2}_{+})]\int_{0}^{t}\mathbb{E}[X^{2}(s)]ds.$ (23)
Applying the Gronwall lemma shows that $\mathbb{E}\left[X^{2}(T_{*})\right]$ is finite; that is, at $t=T_{*}$, $X(T_{*})$ doesn’t explode. Now, if we use $X(T_{*})$ as the initial condition for the next interval $T_{*},2T_{*}]$ and repeat the argument for each interval until $nT_{*}\geq T$, we obtain the global existence of the solution on $[0,T]$, i.e. the solution is in $\mathcal{M_{T}}$.
Step 4. Lastly, we will show that a solution of ($\mathcal{GM}$}) always stays in the class $\mathcal{M_{T}}$ for all $t\in[0,T]$. For integer $n\geq 1$, define the stopping time
$\tau_{n}=T\wedge\inf\{t\geq 0:|X(t)|\geq n\}.$ (24)
Then, $\tau_{n}\uparrow T$ $\mathbb{P}$-a.s. Set $X_{n}(t)=X(t\wedge\tau_{n})$ for $t\in[0,T]$. Then following the same arguments as before provides
$\mathbb{E}[X^{2}_{n}(t)]$ $\leq 3\mathbb{E}[x_{0}^{2}]+3\mu^{2}_{+}\mathbb{E}\left[\left(\int_{0}^{t\wedge\tau_{n}}X_{n}^{p}(s)ds\right)^{2}\right]+3\sigma^{2}_{+}\mathbb{E}\left[\left(\int_{0}^{t\wedge\tau_{n}}X_{n}^{q}(s)dW(s)\right)^{2}\right]$
$\leq 3\mathbb{E}[x_{0}^{2}]+6\mu^{2}_{+}K^{2}(T+1)\mathbb{E}\left[\int_{0}^{t\wedge\tau_{n}}X^{2}_{n}(s)ds\right]+3\sigma^{2}_{+}\mathbb{E}\left[\int_{0}^{t}\mathbf{1}_{[0,\tau_{n}]}(s)X^{2q}_{n}(s)ds\right]$
$\leq 3\mathbb{E}[x_{0}^{2}]+6K^{2}(T+1)(\mu^{2}_{+}+\sigma^{2}_{+})\mathbb{E}\left[\int_{0}^{t\wedge\tau_{n}}X^{2}_{n}(s)ds\right]$
$\leq c_{1}+c_{2}\int_{0}^{t}\mathbb{E}[X^{2}_{n}(s)]ds$ (25)
where $c_{1}=3\mathbb{E}[x_{0}^{2}]$ and $c_{2}=3[2(T+1)K^{2}(\mu^{2}_{+}+\sigma^{2}_{+})]$ are independent of $n$. If we define $f_{n}(t)=\mathbb{E}[X^{2}_{n}(t)]$, and apply the Gronwall lemma to (3) we obtain
$f_{n}(T)\leq c_{1}e^{c_{2}T}=c_{3}\Longrightarrow\sup_{n}f_{n}(T)<\infty\,\,\text{as }n\to\infty.$ (26)
Since $X_{n}(t)\to X(t)$ $\mathbb{P}$-a.s. as $n\to\infty$, by Fatou’s lemma it reads
$\mathbb{E}[X^{2}(t)]=\mathbb{E}[\liminf X^{2}_{n}(t)]\leq\liminf\mathbb{E}[X^{2}_{n}(t)]\leq\sup_{n}\mathbb{E}[X^{2}_{n}(t)],$ (27)
which implies that $\mathbb{E}[X^{2}(t)]$ is bounded and belongs to Lebesgue space $L([0,T])$. Therefore, the solution $X(t)$ of ($\mathcal{GM}$}) remains in the class $\mathcal{M_{T}}$ for any $T$. ∎
Acknowledgments
This work was supported by Athabasca University Research Incentive Account [140111 RIA].
ORCID
https://orcid.org/0000-0002-6001-627X